# references.bib

### v3

@@ 1.1.3. Vanishing bound
The vector spaces $P_{\lambda}$ are finite dimensional, and we can give a rough bound on the cohomological degrees for which they possibly do not vanish. We let $\mathcal{H}_{\lambda}\coloneqq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$, where $d_{\lambda}=\dim V^{\lambda}-\dim G^{\lambda}$.
@@
The cohomologically graded vector space $P_{\lambda}$ given by Theorem 1.1 are concentrated in cohomological degrees $[\dim G^{\lambda}-\dim V^{\lambda},\dim G^{\lambda}]$, that is $\mathcal{P}_{\lambda}^{k}=0$ unless $k\in[\dim G^{\lambda}-\dim V^{\lambda},\dim G^{\lambda}]$.
@@ 2. $\bullet$
    If $X$ is a complex algebraic variety acted upon by an algebraic group $G$, we let $X/G$ be the quotient stack.
@@ 3. $\bullet$
    The multiplicative group is denoted by $\mathbf{G}_{\mathrm{m}}$. The Abelian group of characters of an algebraic torus $T$ is $\mathrm{Hom}_{\mathbf{Z}}(T,\mathbf{G}_{\mathrm{m}})$. The set of cocharacters of $T$ is the Abelian group $\mathrm{Hom}_{\mathbf{Z}}(\mathbf{G}_{\mathrm{m}},T)$.
@@ 4. $\bullet$
    Let $H\subseteq G$ be algebraic groups and $X$ an $H$-variety. We let $X\times^{H}G\coloneqq(X\times G)/H$ where $H$ acts on $X\times G$ by $h\cdot(x,g)=(h\cdot x,gh^{-1})$. We formula $g\cdot(x,g^{\prime})\coloneqq(x,gg^{\prime})$ gives a $G$-action on $X\times^{H}G$.
@@ 6. $\bullet$
    If $G$ is a possibly disconnected algebraic group, then we denote by $G^{0}$ its neutral component.
@@ 1.3. Acknowledgements
At various stages of the preparation of this work, the author was supported by the Royal Society and by the National Science Foundation under Grant No. DMS-1928930 and by the Alfred P. Sloan Foundation under grant G-2021-16778, while the author was in residence at the Simons Laufer Mathematical Sciences Institute (formerly MSRI) in Berkeley, California, during the Spring 2024 semester. The author thanks the SLMath Sciences Institute and the University of Edinburgh for the excellent working conditions. The author is grateful to Ben Davison for the support provided, in particular via postdoctoral fellowships, and to Woonam Lim for pointing out a mistake in the example section.
@@ 2.1. The induction diagram
Let $G$ be a reductive group and $V$ a finite-dimensional representation of $G$. For a cocharacter $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow G$, we let $G^{\lambda}\coloneqq\{g\in G\colon\lambda(t)g\lambda(t)^{-1}=g\}$ be the centraliser of $\lambda$, a Levi subgroup of $G$ and $V^{\lambda}\coloneqq\{v\in V\mid\lambda(t)\cdot v=v\}$ be the fixed locus, a representation of $G^{\lambda}$. We also let $P_{\lambda}\coloneqq\{g\in G\mid\lim_{t\rightarrow 0}\limits\lambda(t)g\lambda(t)^{-1}\text{ exists}\}$, a parabolic subgroup of $G$ and $V^{\lambda\geqslant 0}\coloneqq\{v\in V\mid\lim_{t\rightarrow 0}\limits\lambda(t)\cdot v\text{ exists}\}$, a representation of $P_{\lambda}$.
@@
We let $\mathcal{W}^{\prime}(V)$ be the set of weights of the representation $V$. We have a direct sum decomposition
@@
where $V_{\alpha}$ consists of 00 and nonzero vectors of weight $\alpha$.
We let $\mathcal{W}(V)$ be the collection of weights of $V$, counted multiplicities: $\alpha\in\mathcal{W}^{\prime}(V)$ appears $\dim V_{\alpha}$ times in $\mathcal{W}(V)$.
@@
between characters and cocharacters. We let $\mathcal{W}^{\lambda>0}(V)\coloneqq\{\alpha\in\mathcal{W}(V)\mid\langle\lambda,\alpha\rangle>0\}$ and we define similarly $\mathcal{W}^{\lambda=0}(V)$, $\mathcal{W}^{\lambda\geqslant 0}(V)$, $\mathcal{W}^{\lambda<0}(V)$ and $\mathcal{W}^{\lambda\leqslant 0}(V)$ and their versions, by forgetting the multiplicities.
@@
We consider the commutative induction diagram
[…]
where $V/\!\!/G\coloneqq\mathrm{Spec}(\mathbf{C}[V]^{G})$ is the affine GIT quotient of $V$ by $G$.
@@
$\widetilde{p}_{\lambda}\colon\widetilde{V}^{\lambda\geqslant 0}\coloneqq V^{\lambda\geqslant 0}\times^{P_{\lambda}}G\rightarrow V,$
which is projective. Therefore, $p_{\lambda}$ is representable and projective.
The fact that $q_{\lambda}$ is smooth, surjective of relative dimension #$\#\mathcal{W}^{\lambda>0}(V)-\mathcal{W}^{\lambda>0}(\mathfrak{g})$ is easily seen. That it is in addition a vector bundle stack is a rather classical fact (which is not used in this paper). It comes from the facts that $V^{\geqslant\lambda}\rightarrow V^{\lambda}$ is a vector bundle (by the Białinicky-Birula decomposition [bialynicki1973some]) and the map $P_{\lambda}\rightarrow G^{\lambda}$ has kernel the unipotent radical of $P_{\lambda}$. ∎
@@
The arguments are parallel to that of the proof [meinhardt2019donaldson, Lemma 2.1]. Namely, we consider the commutative diagram
[…]
where $\widetilde{V}^{\lambda\geqslant 0}\coloneqq V^{\lambda\geqslant 0}\times^{P_{\lambda}}G$, so that we have an equivalence of stacks $V^{\lambda\geqslant 0}/P_{\lambda}\simeq\widetilde{V}^{\lambda\geqslant 0}/G$. We now explain the definitions of the maps in this diagram.
The map $q_{\lambda}$ is given by the equivariant map $(V^{\lambda\geqslant 0},P_{\lambda})\rightarrow(V^{\lambda},G^{\lambda})$
Therefore, the pullback of functions by $q_{\lambda}$ induces a morphism
@@
$\overline{q}_{\lambda}\colon\mathrm{Spec}(\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]^{G^{\lambda}})\rightarrow\mathrm{Spec}(\mathbf{C}[V^{\lambda}])=V^{\lambda}/\!\!/G^{\lambda}.$
The map $\jmath_{\lambda}$ is given by the equivariant morphism $(V^{\lambda},G^{\lambda})\rightarrow(V^{\lambda\geqslant 0},P_{\lambda})$ which induces the morphism
@@
Since $q_{\lambda}\circ\jmath_{\lambda}=\mathrm{id}$, we have $\overline{q}_{\lambda}\circ\overline{\jmath}_{\lambda}=\mathrm{id}$. On the ring of functions, we have $\overline{\jmath}_{\lambda}^{*}\circ\overline{q}_{\lambda}^{*}=\mathrm{id}$ and so $\jmath_{\lambda}^{*}$ is surjective. This tells us that $\overline{\jmath}_{\lambda}$ is a closed immersion. Moreover, $\imath_{\lambda}=\overline{p}_{\lambda}\circ\overline{\jmath}_{\lambda}$ and so it suffices to prove that $\overline{p}_{\lambda}$ is a finite map. Since it is a morphism of finite type complex schemes, it suffices to prove that it is an integral morphism [stacks-project, Lemma 01WJ].
Since $\widetilde{p}_{\lambda}\colon\widetilde{V}^{\lambda\geqslant 0}\rightarrow V$ is a projective morphism (Proof of Lemma 2.1), $(\widetilde{p}_{\lambda})_{*}\mathcal{O}_{\widetilde{V}^{\lambda\geqslant 0}}$ is a coherent $\mathcal{O}_{V}$-module. Therefore, the map $\mathbf{C}[V]\rightarrow\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]$ is finite and hence integral [stacks-project, Lemma 01WJ]. We deduce that the map
@@
is integral, and hence finite, as follows. Let $a\in\mathbf{C}[\widetilde{V}^{\geqslant\lambda}]^{G}$. It is solution of a monic polynomial with coefficients in $\mathbf{C}[V]$. By applying the Reynolds operator to the equation, we obtain a monic equation with coefficients in $\mathbf{C}[V]^{G}$ of which $a$ is a zero. This concludes.
For $\lambda\in X_{*}(T)$, we define $d_{\lambda}\coloneqq\#\mathcal{W}^{\lambda=0}(V)-\#\mathcal{W}^{\lambda=0}(\mathfrak{g})=\dim V^{\lambda}/G^{\lambda}$. The quantity $r_{\lambda}$ is defined in Lemma 2.1.
@@
We say that a representation $V$ of a reductive group $G$ is symmetric if $V$ and its dual $V^{*}$ have the same sets of weights counted with multiplicities: $\mathcal{W}(V)=\mathcal{W}(V^{*})$.
For later use, we give the following easy lemma.
@@
Let $V$ be a symmetric representation of $G$. Then,
@@
and by symmetry, #$\#\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\lambda<0}(V)$. Therefore, #$\#\mathcal{W}(V)=2\#\mathcal{W}^{\lambda>0}(V)+\#\mathcal{W}^{\lambda=0}(V)$. These equalities are also valid for $V=\mathfrak{g}$ under the adjoint action of $G$. We obtain the lemma by combining them together. ∎

### v6

@@ 1.3.3. Vanishing bound
The vector spaces $\mathcal{P}_{\lambda}$ given by Theorem 1.1 are finite dimensional (Proposition 4.16), and we can give a rough bound on the cohomological degrees for which they possibly do not vanish. We let $\mathcal{H}_{\lambda}\coloneqq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$, where $d_{\lambda}=\dim V^{\lambda}-\dim G^{\lambda}$. The natural grading on $\mathcal{H}_{\lambda}$ is given by $\mathcal{H}_{\lambda}^{k}=\mathrm{H}^{k+d_{\lambda}}(V^{\lambda}/G^{\lambda})$. It induces a cohomological grading on $\mathcal{P}_{\lambda}$.
@@
The cohomologically graded vector space $\mathcal{P}_{\lambda}$ given by Theorem 1.1 are concentrated in cohomological degrees $\dim G^{\lambda}-\dim V^{\lambda},\dim G^{\lambda}]$, that is $\mathcal{P}_{\lambda}^{k}=0$ unless $k\in[\dim G^{\lambda}-\dim V^{\lambda},\dim G^{\lambda}]$.
@@ 2. $\bullet$
    If $X$ is a complex algebraic variety acted upon by an algebraic group $G$, we let $X/G=[X/G]$ be the quotient stack, that is we omit the brackets.
@@ 3. $\bullet$
    The multiplicative group is denoted by $\mathbf{G}_{\mathrm{m}}$. The Abelian group of characters of an algebraic torus $T$ is $\mathrm{Hom}_{\mathbf{Z}}(T,\mathbf{G}_{\mathrm{m}})$. The set of cocharacters of $T$ is the Abelian group $\mathrm{Hom}_{\mathbf{Z}}(\mathbf{G}_{\mathrm{m}},T)$. We have natural embedding $X_{*}(T)\subseteq\mathfrak{t}\coloneqq\mathrm{Lie}(T)$ and $X^{*}(T)\subseteq\mathfrak{t}^{*}$ respectively given by $\lambda\mapsto\frac{\mathrm{d\lambda}}{\mathrm{d}t}_{|t=1}$ and $\alpha\mapsto\mathrm{d}\alpha(1)$, the differential of a character $\alpha$ at the neutral element $1\in T$. Moreover, these inclusions are compatible with the natural pairings $\langle-,-\rangle\colon X_{*}(T)\times X^{*}(T)\rightarrow\mathbf{Z}$ and $\mathfrak{t}\times\mathfrak{t}^{*}\rightarrow\mathbf{C}$.
@@ 4. $\bullet$
    If $A$ is an Abelian group and $S\subseteq A$ is a subset of $A$, we let $S\coloneqq\{-s\colon s\in A\}$ be the opposite of $S$. We will only consider the case $A=X^{*}(T)$.
  5. $\bullet$
    Let $H\subseteq G$ be algebraic groups and $X$ an $H$-variety. We let $X\times^{H}G\coloneqq(X\times G)/H$ where $H$ acts on $X\times G$ by $h\cdot(x,g)=(h\cdot x,gh^{-1})$. The formula $g\cdot(x,g^{\prime})\coloneqq(x,gg^{\prime})$ gives a $G$-action on $X\times^{H}G$.
@@ 7. $\bullet$
    If $G$ is a possibly disconnected algebraic group, then we denote by $G^{\circ}$ its neutral component.
  8. $\bullet$
    Given a scheme $X$, we denote by $\mathcal{O}_{X}$ its sheaf of regular functions. If $X$ is a complex scheme, then $\mathbf{C}[X]\coloneqq\mathcal{O}_{X}(X)$ set ring of regular functions on $X$.
  9. $\bullet$
    If $S$ is a set (for example, a vector space) with the action of a group $G$, we denote by $S^{G}$ the fixed subset.
@@ 1.5. Acknowledgements
At various stages of the preparation of this work, the author was supported by the Royal Society, by the National Science Foundation under Grant No. DMS-1928930 and by the Alfred P. Sloan Foundation under grant G-2021-16778, while the author was in residence at the Simons Laufer Mathematical Sciences Institute (formerly MSRI) in Berkeley, California, during the Spring 2024 semester. The author thanks the SLMath Sciences Institute and the University of Edinburgh for the excellent working conditions. The author is grateful to Ben Davison for the support provided, in particular via postdoctoral fellowships, to Woonam Lim for pointing out a mistake in the example section, to Dragoş Frăţilă for interesting comments, to Sebastian Schlegel Mejia for suggesting improvements and corrections, to Dimitri Wyss for interesting related discussions, and to Andrei Neguţ for suggesting me to investigate weakenings of the symmetricity condition, which led to the current version involving weakly symmetric representations (Definition 2.3). I am grateful to Daniel Juteau for discussions on perverse sheaves. I would also like to thank the anonymous referee for allowing me to improve this paper.
@@ 2.1. The induction diagram
Let $G$ be a reductive group, $T\subseteq G$ a maximal torus of $G$ and $V$ a finite-dimensional representation of $G$. For a cocharacter $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow T$, we let $G^{\lambda}\coloneqq\{g\in G\colon\lambda(t)g\lambda(t)^{-1}=g\}$ be the centraliser of $\lambda$, a Levi subgroup of $G$ [DM20a, Proposition 1.22] and $V^{\lambda}\coloneqq\{v\in V\mid\lambda(t)\cdot v=v\}$ be the fixed locus, a representation of $G^{\lambda}$. We also let $P_{\lambda}=G^{\lambda\geqslant 0}\coloneqq\{g\in G\mid\lim_{t\rightarrow 0}\limits\lambda(t)g\lambda(t)^{-1}\text{ exists}\}$, a parabolic subgroup of $G$ and $V^{\lambda\geqslant 0}\coloneqq\{v\in V\mid\lim_{t\rightarrow 0}\limits\lambda(t)\cdot v\text{ exists}\}$, a representation of $P_{\lambda}$.
@@
We let $\mathcal{W}^{\prime}(V)\subseteq X^{*}(T)$ be the set of weights of the representation $V$. We have a direct sum decomposition
@@
where $V_{\alpha}$ consists of 0 and nonzero vectors of weight $\alpha$, where $V_{\alpha}\coloneqq\{v\in V\mid t\cdot v=\alpha(t)v\text{ for any $t\in T$}\}$.
We let $\mathcal{W}(V)$ be the collection of weights of $V$, counted with multiplicities: $\alpha\in\mathcal{W}^{\prime}(V)$ appears $\dim V_{\alpha}$ times in $\mathcal{W}(V)$.
@@
between characters and cocharacters. If $(\lambda,\alpha)\in X_{*}(T)\times X^{*}(T)$, then $\alpha\circ\lambda\in\mathrm{End}(\mathbf{G}_{\mathrm{m}})$, and so it is given by raising to some power, which is $\langle\lambda,\alpha\rangle\in\mathbf{Z}$. We let $\mathcal{W}^{\lambda>0}(V)\coloneqq\{\alpha\in\mathcal{W}(V)\mid\langle\lambda,\alpha\rangle>0\}$ and we define similarly $\mathcal{W}^{\lambda=0}(V)$, $\mathcal{W}^{\lambda\geqslant 0}(V)$, $\mathcal{W}^{\lambda<0}(V)$ and $\mathcal{W}^{\lambda\leqslant 0}(V)$ and their versions, by forgetting the multiplicities.
@@
We consider the (commutative) induction diagram
(2.1) $V^{\lambda\geqslant 0}/P_{\lambda}$$V^{\lambda}/G^{\lambda}$${{V^{\lambda}/\!\!/G^{\lambda}}}V//G{{V/\!\!/G}}$$\scriptstyle{q_{\lambda}}$$\scriptstyle{p_{\lambda}}$$\scriptstyle{\pi_{\lambda}}$$\scriptstyle{\pi}$$\scriptstyle{\imath_{\lambda}}$
where $V/\!\!/G\coloneqq\mathrm{Spec}(\mathbf{C}[V]^{G})$ is the affine GIT quotient of $V$ by $G$ and the map $\imath_{\lambda}$ is induced by the equivariant closed immersion $(V^{\lambda},G^{\lambda})\rightarrow(V,G)$.
@@
$\widetilde{p}_{\lambda}\colon\widetilde{V}^{\lambda\geqslant 0}\coloneqq V^{\lambda\geqslant 0}\times^{P_{\lambda}}G\rightarrow V,\quad(v,g)\mapsto g\cdot v\,$,
which can be factored as a closed immersion followed by a projective map
$\begin{matrix}V^{\lambda\geqslant 0}\times^{P_{\lambda}}G&\rightarrow&V\times G/P_{\lambda}&\rightarrow&V\\ (v,g)&\mapsto&(g\cdot v,gP_{\lambda})&&\\ &&(v,gP_{\lambda})&\mapsto&v\end{matrix}$
and is therefore projective. Therefore, $p_{\lambda}$ is representable and projective.
The fact that $q_{\lambda}$ is smooth, surjective of relative dimension $\mathcal{W}^{\lambda>0}(V)-\mathcal{W}^{\lambda>0}(\mathfrak{g})$ is easily seen. That it is in addition a vector bundle stack is a rather classical fact (which is not used in this paper). It comes from the facts that $V^{\lambda\geqslant 0}\rightarrow V^{\lambda}$ is a vector bundle (by the Białinicky-Birula decomposition [Bia73], but this be seen directly in the case of linear actions) and the kernel of the map $P_{\lambda}\rightarrow G^{\lambda}$ is the unipotent radical of $P_{\lambda}$. ∎
@@
The arguments are parallel to that of the proof [MR19, Lemma 2.1]. Namely, we consider the commutative diagram
$V^{\lambda\geqslant 0}/P_{\lambda}$$V^{\lambda}/G^{\lambda}$$\mathrm{Spec}(\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]^{G})$${{V^{\lambda}/\!\!/G^{\lambda}}}V//G{{V/\!\!/G}}$$\scriptstyle{q_{\lambda}}$$\scriptstyle{\jmath_{\lambda}}$$\scriptstyle{\pi_{\lambda\geqslant 0}}$$\scriptstyle{p_{\lambda}}$$\scriptstyle{\pi_{\lambda}}$$\scriptstyle{\overline{q}_{\lambda}}$$\scriptstyle{\overline{p}_{\lambda}}$$\scriptstyle{\pi}$$\scriptstyle{\overline{\jmath}_{\lambda}}$$\scriptstyle{\imath_{\lambda}}$
where $\widetilde{V}^{\lambda\geqslant 0}\coloneqq V^{\lambda\geqslant 0}\times^{P_{\lambda}}G$ (see §1.4), so that we have an equivalence of stacks $V^{\lambda\geqslant 0}/P_{\lambda}\simeq\widetilde{V}^{\lambda\geqslant 0}/G$. We now explain the definitions of the maps in this diagram.
The map $q_{\lambda}$ is given by the equivariant map $(V^{\lambda\geqslant 0},P_{\lambda})\rightarrow(V^{\lambda},G^{\lambda})$, $(v,g)\mapsto(\lim_{t\rightarrow 0}\limits\lambda(t)\cdot v,\lim_{t\rightarrow 0}\limits\lambda(t)g\lambda(t)^{-1})$. Therefore, the pullback of functions by $q_{\lambda}$ induces a morphism between rings of invariant functions
@@
$\overline{q}_{\lambda}\colon\mathrm{Spec}(\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]^{G})\rightarrow\mathrm{Spec}(\mathbf{C}[V^{\lambda}]^{G^{\lambda}})=V^{\lambda}/\!\!/G^{\lambda}\,.$
The map $\jmath_{\lambda}$ is given by the equivariant morphism $(V^{\lambda},G^{\lambda})\rightarrow(V^{\lambda\geqslant 0},P_{\lambda})$, $(v,g)\mapsto(v,g)$ which induces the morphism of algebras
@@
Since $q_{\lambda}\circ\jmath_{\lambda}=\mathrm{id}$, we have $\overline{q}_{\lambda}\circ\overline{\jmath}_{\lambda}=\mathrm{id}$. On the ring of functions, we have $\overline{\jmath}_{\lambda}^{*}\circ\overline{q}_{\lambda}^{*}=\mathrm{id}$ and so $\overline{\jmath}_{\lambda}^{*}$ is surjective. This tells us that $\overline{\jmath}_{\lambda}$ is a closed immersion. Moreover, $\imath_{\lambda}=\overline{p}_{\lambda}\circ\overline{\jmath}_{\lambda}$ and so it suffices to prove that $\overline{p}_{\lambda}$ is a finite map. Since it is a morphism of finite type complex schemes, it suffices to prove that it is an integral morphism [Sta24, Lemma 01WJ].
Since $\widetilde{p}_{\lambda}\colon\widetilde{V}^{\lambda\geqslant 0}\rightarrow V$ is a projective morphism (Proof of Lemma 2.1), $(\widetilde{p}_{\lambda})_{*}\mathcal{O}_{\widetilde{V}^{\lambda\geqslant 0}}$ is a coherent $\mathcal{O}_{V}$-module. Therefore, the map $\mathbf{C}[V]\rightarrow\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]$ is finite and hence integral [Sta24, Lemma 01WJ]. We deduce that the map
@@
is integral, and hence finite, as follows. Let $a\in\mathbf{C}[\widetilde{V}^{\geqslant\lambda}]^{G}$. It is solution of a monic polynomial with coefficients in $\mathbf{C}[V]$. By applying the Reynolds operator to the equation, we obtain a monic equation with coefficients in $\mathbf{C}[V]^{G}$ of which $a$ is a zero. This concludes the proof.
We define the equivalence relation $\equiv$ on $X_{*}(T)$ by $\alpha\equiv\beta\iff\exists a,b\in\mathbf{Z}_{>0},\quad a\alpha=b\beta$. The quotient set is denoted by $\overline{X_{*}(T)}\coloneqq X_{*}(T)/\equiv$. If $S\subseteq X_{*}(T)$, we denote by $\overline{S}$ its image in $\overline{X_{*}(T)}$. Similarly, if $S$ is a collection of characters (with multiplicities allowed), we denote by $\overline{S}$ the corresponding collection of elements of $\overline{X_{*}(T)}$ (with multiplicities taken into account).
@@
We say that a representation $V$ of a reductive group $G$ is
  1. (1)
    symmetric if $V$ and its dual $V^{*}$ have the same sets of weights counted with multiplicities: $\mathcal{W}(V)=\mathcal{W}(V^{*})$.
  2. (2)
    weakly symmetric if $\overline{\mathcal{W}(V)}=\overline{\mathcal{W}(V^{*})}$.
Since a representation of a reductive group is determined by its character, a representation $V$ of $G$ is symmetric if and only if it is self-dual, that is $V\cong V^{*}$ as representations of $G$.
Example 2.4.
The following are symmetric representations:
  1. (1)
    The trivial representation $\{0\}$ of a reductive group $G$,
  2. (2)
    The adjoint representation $\mathfrak{g}=\mathrm{Lie}(G)$ of a reductive group $G$,
  3. (3)
    The cotangent representation $\mathrm{T}^{*}V$ of any representation $V$ of a reductive group $G$,
  4. (4)
    The representation space of $\mathbf{d}\in\mathbf{C}^{Q_{0}}$-dimensional representations of a symmetric quiver $Q=(Q_{0},Q_{1})$, for $G=\prod_{i\in Q_{0}}\mathrm{GL}_{\mathbf{d}_{i}}$.
  5. (5)
    Any representations of products of simple groups of types $B_{n},C_{n},D_{n}\text{ ($n$ even)},E_{7},E_{8},F_{4},G_{2}$. Indeed, for such a group $G$, $w_{0}=-1\in W$ and for any $w\in W$, $\dim V_{\alpha}=\dim V_{w\cdot\alpha}$ [BB07, Planches I–IX], [Hum90, Exercise and Table 1 page 59].
Example 2.5.
The two-dimensional representation of $\mathbf{C}^{*}$ given by $t\cdot(u,v)=(tu,t^{-2}v)$ is weakly symmetric but not symmetric.
For $\lambda\in X_{*}(T)$, we define $d_{\lambda}\coloneqq\#\mathcal{W}^{\lambda=0}(V)-\#\mathcal{W}^{\lambda=0}(\mathfrak{g})=\dim V^{\lambda}/G^{\lambda}$. The quantity $r_{\lambda}$ is defined in Lemma 2.1. For later use, we give the following easy lemma.
@@
Let $V$ be a weakly symmetric representation of $G$. Then,
@@
and by weak symmetry (Definition 2.3), $\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\lambda<0}(V)$. Therefore, $\mathcal{W}(V)=2\#\mathcal{W}^{\lambda>0}(V)+\#\mathcal{W}^{\lambda=0}(V)$. These equalities are also valid for $V=\mathfrak{g}$ under the adjoint action of $G$ (which is a symmetric representation, Example 2.4). We obtain the lemma by combining them together. ∎